OpenStudy (anonymous):

Can someone help me out with this. I dont want the answer. i just want help figuring it out. I'll post it below.

OpenStudy (anonymous):

\[\frac{ 2a^2+a-6 }{ a^2+4a-5 }\times \frac{ a^3-3a^2+2a }{ 4a^2-6a }\]

OpenStudy (phi):

This is an exercise in factoring the polynomials. can you factor each one? (or any of them!?)

OpenStudy (anonymous):

Yes

OpenStudy (phi):

the "easy" way to multiply is multiply top*top and bottom * bottom it's clear we will get very ugly polynomial divided by very ugly polynomial however, if these factor (and they probably do), the same factor in the bottom and the top will cancel. After we cancel all that we can, we multiply out what is left ... probably something relatively simple

OpenStudy (anonymous):

thats the problem where i'm at is factoring them.

OpenStudy (phi):

which ones can you factor?

OpenStudy (anonymous):

i got one of bottom ones

OpenStudy (phi):

can you post what you got?

OpenStudy (anonymous):

its 2a(2a-3)

OpenStudy (phi):

ok. now let's try the other bottom polynomial a^2 +4a - 5 1) look at the last number and list all pairs that multiply to give you 5 1,5 is the only pair 2) note the sign on the -5. that means the "roots' will have *different* signs 3) look at the +4 on 4a the + sign means the biggest root is positive because we have different signs, the smaller root is negative 4) put in the proper signs on the pairs listed in step 1) -1+5 5) add each pair (in this case we only have -1+5) +4 6) if the sum matches the +4 in 4a then we found our factors in this case -1 and 5 write (x-1)(x+5)

OpenStudy (phi):

let's try a^3 -3a^2+2a first, factor a from each term. what do you get ?

OpenStudy (anonymous):

im not very sure with the a^3 and -3^2

OpenStudy (phi):

remember the little 3 in the upper right is a short cut to show a*a*a so we could write it as a*a*a - 3*a*a +2*a can you factor a from each term?

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get?

OpenStudy (anonymous):

I'm not entirely sure.

OpenStudy (phi):

example of factoring: if you have a*a + a you can divide each term by a: a*a/a = a and a/a = 1 to get (a+1) multiplied by the a we "factored out" a(a+1) if we distribute the a (multiply a times each term inside the parens) we get a^2 +a showing that a(a+1) is the same as a^2 +a

OpenStudy (phi):

a*a*a - 3*a*a +2*a can you factor a from each term?

OpenStudy (phi):

we "take" an "a" out of each term (technically we divide) and write the a outside the parens

OpenStudy (anonymous):

would it be a or a^2 outside the ( )

OpenStudy (phi):

you can factor out something that is *shared* by all the terms if you have a^3 +a^2 which means a*a*a + a*a both terms have a*a , so we can factor out a*a

OpenStudy (phi):

but notice in a^3 -3a^2+2a which means a*a*a -3*a*a + 2*a the last term 2*a has only 1 a. Each term has 1 a, but not 2 a's so we take out an "a" from each term (not a*a, just a)

OpenStudy (anonymous):

would there be two ( ) or one?

OpenStudy (phi):

just one a can be factored out.

OpenStudy (anonymous):

would it be a(a^2-3a+2)

OpenStudy (phi):

yes

OpenStudy (phi):

now we can do more. Factor a^2-3a+2

OpenStudy (anonymous):

\[\frac{ (2a-3)(a+2) }{ (a-1)(a+5) }\times \frac{ a(a^2-3a+2) }{ 2a(2a-3) }\] is this what i should have right now

OpenStudy (phi):

yes that looks good but you can factor a^2 -3a + 2 can you use the steps I posted up above?